OpenStudy (anonymous):

Factor the trinomial y^3-11^2+28y

OpenStudy (anonymous):

y^3 - 7y^2 - 4y^2 + 28y

OpenStudy (marco26):

is it really 11^2 or 11y^2?

OpenStudy (anonymous):

Factor the trinomial of 11y^3-11y^2+28y

OpenStudy (marco26):

Have you tried synthetic division using trial and error?

OpenStudy (anonymous):

Check your question again .. something is wrong , Are you sure of 11y^3 ?

OpenStudy (anonymous):

oops sorry I meant y^3-11y^3+28y

OpenStudy (anonymous):

To factor this completely, let's first look at what each of the terms have in common. You'll notice that each term has a y in common. So, we can start by factoring the y. This gives us: y(y^2-11y+28) Next, we need to see if we can factor what's in the parenthesis. To do so, we need to figure out if there are two numbers, that when multiplied together, give us 28, and when added together, give us -11. We know that both numbers must be negative, because 2 negatives multiplied together give us a positive, and 2 negative numbers added together, will give us a negative. Let's find these two numbers: -1 x -28 = 28; -1 + -28 = -29 WILL NOT WORK -2 x -14 = 28; -2 + -14 = -16 WILL NOT WORK -4 x -7 = 28; -4 + -7 = -11 WORKS!!! Now that we know -4 and -7 works, we will put these two numbers into factor form: (y - 4)(y - 7) Tack on the y that we factored out originally, to the beginning of these factors, and we will have our final answer: y(y - 4)(y - 7) y=0 y-4=0 , y=4 y-7=0 , y=7 ======================== You got it ?

OpenStudy (anonymous):

Yes I do understand it. thank you very much Eyad! this was helpful

OpenStudy (anonymous):

Yw :).

OpenStudy (anonymous):

so my solution I can write it out like this (y-4)(y-7)

OpenStudy (anonymous):

It's a prime right?